OpenStudy (anonymous):

I need help understanding this: If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 5x - 2y - z = 16 x + y + 4z = 6 3x + 2y + 6z = 16 Thanks!

OpenStudy (cwrw238):

from second equation x = 6 - y - 4z now plug this for x into equations 1 and 3

OpenStudy (anonymous):

So, something like 5( 6-y-4z)=16?

OpenStudy (anonymous):

and 3( 6-Y-4Z) + 2y + 6z = 16? And then would I distribute those?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

for the first one its 5(6 - y - 4z) -2y - z = 16

OpenStudy (anonymous):

So for the first one I would get 30 - 5y - 20z + 16???

OpenStudy (anonymous):

oh

OpenStudy (cwrw238):

no you missed out the -2y - z

OpenStudy (anonymous):

So, 30 - 5y - 20z - 2y -z = 16?

OpenStudy (cwrw238):

yes and that simplifies some more to 30 - 7y - 21z - 16 = 0 14 - 7y - 21z = 0

OpenStudy (anonymous):

My options are a) -3y +19z = -14 5y +18z =-2 b) 7x +7z =28 x-2z=4 c) -7y-21z=-14 -y-6z=-2 d) 21x - 7y=70 33x-10y=112 I got the simplified version you wrote but how do I get to one of the answers

OpenStudy (cwrw238):

or if you prefer it with a number on RHS: 7y + 21z = 14

OpenStudy (cwrw238):

right well my equation is really the same as the first one in c)

OpenStudy (anonymous):

Okay and the second part of the equation in c is just a simplified version?

OpenStudy (cwrw238):

i'm checking that one now 3( 6-Y-4Z) + 2y + 6z = 16? 18 -3y - 12z +2y + 6z = 16 18 - y - 6z = 16 -y - 6z = -2 yup c) is the correct option

OpenStudy (anonymous):

Thanks so much :)!

OpenStudy (cwrw238):

yw